EZ005

PI: Zhang
Title: AO 0235+164

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: B
Obs time (UT): EU 10:00(18/02)-22:00(18/02)
Stations: EbHhJbMcNtOnSmShTrUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

HartRAO (Jonathan Quick at Tue Feb 18 17:34:29 EST 1997) Minor failures
Read heads failing - write heads still seem OK

Medicina (Marco Bondi at Tue Feb 18 17:50:53 EST 1997) Success
Looks to be ok

Onsala (A. Polatidis at Tue Feb 18 23:11:07 EST 1997) Success
Due to high winds we could only observe for the last 2.5 hours. Data start at UT 1933 (pass 16). Rain and occasionally snow, winds 9-13 m/s. No known problems for the recorded data.

Effelsberg (dgraham at Wed Feb 19 06:05:33 EST 1997) Minor failures
Low elevation at start, underestimated drive time to calibration source.

Torun (Roman Feiler at Wed Feb 19 20:30:59 EST 1997) Success
No known problems. Bad weather; wet snow and wind 5-7 m/s.

Westerbork (Tony Foley at Thu Feb 20 05:23:22 EST 1997) Minor failures
First 2 scans were outside our Hour Angle range, so WSRT started observing at 10:28UT. Strong gusting winds gave some telescope pointing problems ~12-13UT.

Jodrell (Alastair Gunn at Thu Feb 20 12:45:39 EST 1997) Failure
Did not observe due to high winds all day.

Shanghai (Huang Xinyong at Wed Mar 5 04:48:24 EST 1997) Success
Clock offset = 0.05 msec. No problems, weather fine.

Calibration (Fredrik T Rantakyro at Thu May 8 11:32:23 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name ez5.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Scott Aaron at Fri Feb 6 11:42:42 MET 1998) Minor failures
Correlation completed. The third IF at Medicina is bad, so the data were not correlated. Frequent tracking problems at Torun.